# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

### v1

@@ 4. Experiment and Finding
To deeply understand the relationships between models and datasets, we aim to answer seven critical research questions (RQs). We believe they can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development and deployment.
  • RQ #1: What are the properties of the LLM supply chain graph?
  • RQ #2: What structural patterns emerge from the LLM ecosystem?
  • RQ #3: What are the supply chain relationships between datasets?
  • RQ #4: What are the relationships between models and datasets?
  • RQ #5: What insights can be gained as the graph temporally updating?
@@ 4.1. RQ #1: Supply Chain Graph Properties
Figure 3. (a) Indegree distribution and (b) Outdegree distribution. The X-axis represents the number of indegrees and outdegrees, respectively, and the Y-axis represents the number of vertices in a logarithmic scale.
This research question aims to understand the critical properties of the LLM supply chain graph, i.e., graph basics and degree distribution.
Graph basics. The collected supply chain graph is a large directed heterogeneous graph with 397,376 nodes and 453,469 edges as of June 30th, 2025. In particular, there are six different types of nodes, including 25,717 base models, 115,103 fine-tuned, 79,240 adapters, 98,125 quantization models, 13,027 merges and 66,164 datasets. The average degree is about 1.14, denoting that it is a very sparse graph.
The degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph. Figure 3(a) and (b) illustrate the indegree and outdegree distribution of the graph, respectively. We show not only the total distribution but also the distribution of five types of nodes, including base models, fine-tuned models, adapter models, quantization models, and datasets. We made two interesting observations.
(i) This degree distribution in our supply chain graph shows a heavy-tailed behavior. In particular, the indegree distribution shows a large spread across different categories. The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution). The power-law behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph. In particular, the model images from macrocosm-os has the highest indegree value of 550, and Mistral-7B-v0.1 from mistralai has the highest outdegree value of 1,093. Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models. To this end, we can conclude that this graph is a power-law graph, also known as a scale-free graph. That is, a graph whose degree distribution follows a power law distribution. In other words, a small number of nodes have many connections, while most nodes have a few connections.
(ii) We observe that all the base models have at least one outdegree as they are used by the task-specific models. For the dataset nodes, there are 20,093 datasets with zero indegrees, meaning they are the minimum datasets that can no longer be partitioned.
Finding #1: The LLM supply chain graph is a large, sparse graph that exhibits a power-law degree distribution, where a small number of high-degree base models and datasets serve as central hubs.
@@ 4.2. RQ #2: Supply Chain Structural Analysis
Connectivity analysis in a graph is crucial for understanding how well different nodes are linked and helps identify critical nodes and edges that maintain overall connectivity. This graph is a directed acyclic graph (DAG) because there are no dependencies in a cycle form, which also means there are no strongly connected components. Because of that, we only perform weakly connected components (WCC) analysis, and the total number of WCCs in our supply chain graph is 44,417.
Figure 4. Cumulative distribution function (CDF) of WCC size.
Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution. We made two interesting observations. (i) The largest WCC covers 244,063 nodes, accounting for 61.4% of all the nodes. It reflects the dense interconnections that pervade the ecosystem. This vital element is essential for effective information sharing, resource allocation, and structural support and is the base of the ecosystem. In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2.
(ii) In contrast, the remaining 153,313 WCCs collectively hold 31.4% of the nodes, with most having 1, 2, to 3 nodes. This indicates a fragmented outer edge characterized by specialized models, less frequent datasets, or active experimental projects. The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system. For example, tblard/tf-allocine (French sentiment), distilbert-base-uncased-finetuned-sst-2-english (sentiment analysis), and PolyAI/minds14 (banking queries) remain disconnected due to limited reuse or insufficient metadata.
@@
Community analysis is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph. In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets. We use the Louvain method (De Meo et al., 2011), which is a greedy optimization method for maximizing modularity. Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them. It compares the actual density of edges inside communities to the expected density if edges were placed at random.
@@
Table 2 summarizes the top-10 communities detected using the Louvain method. Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.97, indicative of strong intra-community connectivity. Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem.
We made two interesting observations. (i) The largest community (ID 1) consists of 9,158 nodes and attains a modularity score of 0.97, indicating extremely cohesive internal structure. It revolves around base models such as OLMoE and CausalLM, and general-purpose datasets like prompt-perfect. This suggests a densely connected nucleus facilitating widespread reuse and fine-tuning—effectively functioning as the backbone of the LLM ecosystem.
(ii) Several other communities reflect clear task-based segmentation. For instance, Community 3 ( 7,058 nodes, modularity 0.97) focuses on benchmarking, with models like Wanxiang and datasets such as marco-o1. Similarly, Community 6 (5,030 nodes, modularity 0.97) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
Table 2. Top-10 Louvain communities in the HuggingGraph.
@@
1 9158 OLMoE, CausalLM prompt-perfect 0.97
2 7058 qwen2.5_math Marco-o1 0.97
3 6987 Wanxiang smartllama3.1 0.97
4 5696 tinyllama Llama-1B 0.97
5 5355 Qwen2.5-32B Matter-0.2 0.97
6 5030 MedLlama-3-8B dpo-mix 0.97
7 4297 Electra, ArliAI MixEval 0.97
8 4133 aesqwen1.5b llava 0.97
9 4067 bert TORGO 0.97
10 3932 Mistral vicuna_format 0.97
Finding #2: The LLM supply chain graph reveals a structure with a densely connected core and a fragmented periphery. A single dominant component encompasses the majority of nodes, supporting efficient model reuse and dependency tracing, while many smaller, isolated components reflect specialized or under-integrated efforts. Community analysis shows strong modularity, with nodes clustering around shared tasks and functional roles.
4.3. RQ #3: Dataset Supply Chain Analysis
This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph. There are two relationships between datasets. On one hand, one dataset can be created by combining a variety of other datasets from disparate sources. On the other hand, one dataset can serve as a building block for new datasets, which enables continuous improvement in machine learning assets. In the following, we discuss the impacts on both training and benchmarking datasets.
Training dataset impact. In the LLM supply chain graph, 66,164 datasets act as training datasets that are used for model training or generating new datasets. Table 3 presents the top 10 datasets ranked by the highest number of included and derived datasets.
(i) We observe that a training dataset can include many small datasets. Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm. It includes 550 sub-datasets and is a general-purpose image modification corpus used in advance multimodal AI research and large-scale vision-language model training. Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs includes 215 sub-datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
(ii) The other observation is that a single dataset can be included in many larger datasets, showing the huge overlaps between the datasets and also the models trained with them. The right two columns of Table 3 show the top 10 datasets sorted by the maximum number of derived datasets they have been included. The leading contributor, HuggingFaceH4 from HuggingFace has been involved in 989 derived datasets. It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025). Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 derived datasets, is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
Table 3. Top-10 datasets sorted by the # of included and # of derived datasets.
Dataset
# of included
dataset
Dataset
# of derived
datasets
macrocosm-os/images 550 HuggingFaceH4 989
bespokelabs/Bespoke 215 CodeFeedback 857
databricks/databricks 137 MADLAD 850
Open-Orca/OpenOrca 119 Capybara 823
nguha/legalbench 117 Glot500 804
OpenAssistant/oasst1 111 dolphin-coder 756
LDJnr/Capybara 104 SlimOrca 726
kvn420/Tenro_V4.1 101 orca_dpo_pairs 687
teknium/OpenHermes 98 gutenberg 675
Anthropic/hh-rlhf 96 samantha 658
Finding #4: The datasets play critical roles in model training with 66,164 datasets are contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
4.4. RQ #4: Supply Chain between LLM Models and Datasets
This research question aims to explore the interconnections between models and datasets within the supply chain graph. It provides valuable insights from dual perspectives, as discussed below, including one dataset versus multiple models and one model versus multiple datasets.
One dataset versus multiple models refers to the case when a single dataset is used to train multiple models. Table LABEL:tab:dataset_model_1 shows the top 10 datasets based on the number of models trained on them. In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks. It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets. Table LABEL:tab:dataset_model_2 shows the top 10 models ranked by the number of datasets used for training. DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets. Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications. The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets respectively, across fine-tuning and quantized variants (e.g., GGUF format). Interestingly, most of these models are fine-tuned—specifically, 8 out of the top 10—highlighting a trend where base models are transformed into task-specific variants through diverse training datasets. This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
Finding #5: There exists a strong interdependence between models and datasets, where widely adopted datasets like Mistral-7B-v0.1 and TinyLlama-1.1B-v1.0 serve as the base for training hundreds of models across different adaptation techniques, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance their versatility, highlighting the critical role of dataset-model interactions in advancing AI capabilities.
Table 4. Top-10 datasets sorted by # of models trained.
Dataset
@@ 4.4. RQ #4: Supply Chain between LLM Models and Datasets
Merges

### v2

@@ 4. Experiments and Finding
To deeply understand the relationships between models and datasets, we study seven critical research questions (RQs), which can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development.
  • RQ #1: What are the properties of LLM supply chain graph?
  • RQ #2: What structural patterns emerge?
  • RQ #3: What are the supply chain between LLM models?
  • RQ #4: What are the supply chain between datasets?
  • RQ #5: What are the relationships between models and datasets?
  • RQ #6: What insights can be gained from the dynamic updates?
  • RQ #7: How can HuggingGraph be applied to other platforms?
@@ 4.1. RQ #1: Supply Chain Graph Properties
Figure 3. (a) Indegree distribution and (b) Outdegree distribution. X-axis represents degrees, and the Y-axis represents the number of vertices in a logarithmic scale.
This research question aims to understand the critical properties of LLM supply chain graph, i.e., graph basics and degree distribution.
Graph basics. The collected supply chain graph is a medium-scale directed heterogeneous graph with 402,654 nodes and 462,524 edges as of June 30th, 2025. In particular, there are six different types of nodes, including 28,384 base models, 115,211 fine-tuned, 79,254 adapters, 98,143 quantization models, 13,028 merges, and 68,634 datasets. The average degree is about 1.15, denoting that it is a very sparse graph. Furthermore, we identified substantial metadata missing. As of June 30, 2025, among 1.8 million models, only 50,156 (2.76%) include a model tree, while $\sim$ 550K models lack any metadata beyond their name, with nearly 400K entries empty or invalid. Similarly, of the 450K datasets, only 68,634 (15.26%) provide datacards, leaving $\sim$ 380K datasets without any usable metadata.
This highlights a broader issue in the AI community, where a significant number of models and datasets lack consistent and structured documentation on the supply chain. This reflects the need of more transparent disclosure.
Degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph. Figure 3 illustrates the indegree and outdegree distribution of the graph. We show not only the total distribution but also the distribution of six types of nodes, including base models, fine-tuned models, adapter models, quantization models, merged models, and datasets. We made two interesting observations. (i) This degree distribution in our supply chain graph shows a heavy-tailed behavior. In particular, the indegree distribution shows a large spread across different categories. The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution). The heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph. In particular, the dataset macrocosm-os/images has the highest indegree value of 550, and Mistral-7B-v0.1 from “mistral AI” has the highest outdegree value of 1,093. Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models.
Finding #1: The LLM supply chain graph is medium-scale, sparse, and heavy-tailed distribution. However, a significant number of models and datasets lack metadata, highlighting the need for more transparent supply chain documentation.
@@ 4.2. RQ #2: Supply Chain Structural Analysis
Connectivity analysis analyzes the structural connectivity information. We performed weakly connected components (WCC), which identifies maximal sets of nodes that remain connected when edge directions are ignored. The total number of WCCs in our supply chain graph is 44,908. Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution. We made two interesting observations. (i) The largest WCC covers 247,244 nodes, accounting for 61.8% of all the nodes. It reflects the dense interconnections that pervade the ecosystem. This vital element is essential for effective information sharing, resource allocation, and structural support and is the base of the ecosystem. In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2. (ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes (i,e,.155,410 nodes), with most having 1, 2 or 3 nodes. This indicates a fragmented outer edge characterized by specialized models, rare datasets, or active experimental projects. The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system. For example, zhongqy/RMCBench (benchmarking dataset) and yigit69/bert-base-uncased-finetuned-rte-run_3 (recognizing textual entailment task) remain disconnected due to limited reuse or insufficient metadata.
We also computed strongly connected component (SCC), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs. Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 2,169 nodes (0.54% of the graph). The largest SCC comprises 478 dataset nodes, among them tree‑of‑knowledge and OpenHermes‑2.5, forming a tightly‑knit cluster. By contrast, the remaining 99.46% of nodes each reside in trivial (size‑1) SCCs.
@@
Community detection is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph. In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets. We use the Louvain method (De Meo et al., 2011), which is a greedy optimization method for maximizing modularity. Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them.
@@
Table 2 summarizes the top-10 communities detected using the Louvain method. Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.96, indicative of strong intra-community connectivity. Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem. We made two interesting observations. (i) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure. It revolves around base models like OLMoE and CausalLM, and general-purpose datasets like prompt-perfect. This suggests a densely connected cluster facilitating widespread reuse and fine-tuning. (ii) Several other communities reflect clear task-based segmentation. For instance, Community 3 (7,388 nodes, modularity 0.96) focuses on benchmarking, with models like Wanxiang and datasets like marco-o1. Similarly, ID 5 (4,554 nodes, modularity 0.96) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
Table 2. Top-10 Louvain communities sorted by size.
@@
1 9,390 OLMoE, CausalLM prompt-perfect 0.96
2 7,388 qwen2.5_math Marco-o1 0.96
3 6,989 Wanxiang smartllama3.1 0.96
4 6,813 tinyllama Llama-1B 0.96
5 5,163 Qwen2.5-32B Matter-0.2 0.96
6 4,554 MedLlama-3-8B dpo-mix 0.96
7 4,262 Electra, ArliAI MixEval 0.96
8 3,947 aesqwen1.5b llava 0.96
9 3,829 bert TORGO 0.96
10 3,828 Mistral vicuna_format 0.96
Finding #2: The LLM supply chain graph features a dominant core (61.8% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
4.3. RQ #3: Supply Chain Analysis of LLM Models
This research question aims to provide a holistic view of the dependencies between the models within the LLM supply chain, particularly from both base and task-specific models.
Base model impact. We would like to understand the impact of base models. Here, we quantify the impact of a base model as the number of task-specific models that depend on it. The more dependencies, the larger the impact it has. We start with a base model and perform forward analysis by computing a breadth-first search (BFS) following the outgoing edges to get that from our LLM supply chain graph. This leads to a forward subgraph, which denotes all the models that depend on the base model, including fine-tuned, adapted, quantized, or merged models.
Table 3 shows the top-10 base models sorted by the forward subgraph size, which is the number of impacted task-specific models. We make two interesting observations. (i) A base model can significantly impact the LLM supply chain ecosystem. For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024). Due to its relatively small size, which allows for deployment in resource-constrained environments, making advanced AI accessible to broader stakeholders (Grattafiori et al., 2024). It has generated up to 7,544 models, including 1,710 fine-tuned versions, 1,542 adapters, 3,473 quantizations, and 1,693 merged models tailored to specific tasks. (ii) In terms of fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most times, totaling 2,105. It is a faster and lighter version of the Mistral model from Mistral AI. It is obtained by training on a large corpus with grouped-query and sliding window attention by Mistral AI, delivering efficient text generation, NLP, and code assistance on consumer hardware for low-latency applications like chatbots and text classification (Thakkar and Manimaran, 2023).
Table 3. Top-10 base models sorted by forward subgraph size.
Base model Total
Fine-
tune
Adapter
Quanti-
zation
Merge Level
Llama-3.1-8B 7544 1710 1542 3473 1693 25
Mistral-7B-v0.1 6744 2105 2187 1435 1254 27
Qwen2.5-7B 6733 1972 1764 2516 1132 11
Meta-Llama-3-8B 5633 967 1511 2220 1967 21
Llama-3.1-70B 4063 698 281 2075 2519 11
Qwen2.5-32B 3909 1086 158 2311 1049 12
Qwen2.5-1.5B 3645 1300 1290 949 248 8
Qwen2.5-0.5B 3521 1669 1006 810 46 11
Qwen2.5-14B 3362 726 411 1880 1166 15
Meta-Llama-3-8B-Instruct 3118 640 405 1394 1305 34
Task-specific model analysis. Given a task-specific model, we want to understand how it evolves, meaning what other models it has relied on. To achieve that, for each task-specific model, we perform a backward analysis by running a BFS following the incoming edges. To that end, the derived subgraph shows the models it relies on.
Table 4 shows the top 10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on. We make two interesting observations. (i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain. This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model cc4ai, and has undergone extensive lineage evolution before reaching its final form. Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4. (ii) We observed that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
Finding #3: Base models like Llama-3.1-8B dominate the LLM supply chain, spawning thousands of derivatives, while task-specific models such as command-r-1-layer exhibit deep dependencies with other task-specific variants but avoid adapters or merges.
4.4. RQ #4: Dataset Supply Chain Analysis
Table 4. Top-10 models sorted by backward subgraph size.
Model
Model
Type
@@ 4.4. RQ #4: Dataset Supply Chain Analysis
Level Base Model
command-r-1-layer Finetune 40 39 0 39 c4ai
KoModernBERT Finetune 21 20 0 20 ModernBERT
t5-small Finetune 21 20 0 20 t5-small
clinical_260k Finetune 20 19 0 19 clinical_180K
t5-small-finetuned Finetune 17 16 0 16 t5-small
clinical_300k Finetune 16 15 0 15 clinical_180K
clinical_259k Finetune 16 15 0 15 clinical_180K
LeoPARD-0.8.1 Finetune 16 2 13 15 DeepSeek-R1
LeoPARD-0.8.2-4bit Quantization 16 1 14 15 DeepSeek-R1
LeoPARD-0.8.1-4bit Quantization 16 1 14 15 DeepSeek-R1
This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph. There are two relationships between datasets. On one hand, one dataset can be created by combining a variety of other datasets from disparate sources. On the other hand, one dataset can serve as a building block for new datasets.
In the LLM supply chain graph, 68,634 datasets act as training datasets that are used for model training or generating new datasets. Table 5 presents the top 10 datasets ranked by the highest number of included and derived datasets. (i) We observe that a training dataset can include many small datasets. Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm. It is included in 550 datasets and is a general-purpose image modification corpus used in advanced multimodal AI research and large-scale vision-language model training. Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs is part of 215 datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
Table 5. Top-10 datasets by # of included & derived datasets.
Dataset # of included Dataset # of derived
macrocosm-os/images 550 HuggingFaceH4 989
bespokelabs/Bespoke 215 CodeFeedback 857
databricks/databricks 139 MADLAD 850
Open-Orca/OpenOrca 121 Capybara 823
nguha/legalbench 117 Glot500 804
OpenAssistant/oasst1 112 dolphin-coder 756
LDJnr/Capybara 108 SlimOrca 726
kvn420/Tenro_V4.1 105 orca_dpo_pairs 687
teknium/OpenHermes 104 gutenberg 675
Anthropic/hh-rlhf 98 samantha 658
(ii) The other observation is that a single dataset can be derived to many datasets, showing the huge overlaps between the datasets and also the models trained with them. The right two columns of Table 5 show the top 10 datasets sorted by the maximum number of derived datasets they have been included in. The leading contributor, HuggingFaceH4 from HuggingFace has been derived to 989 datasets. It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025). Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 datasets, which is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
Finding #4: Datasets play critical roles in training with 68,634 datasets contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
4.5. RQ #5: Supply Chain between Models and Datasets
This research question aims to explore the interconnections between models and datasets within the supply chain graph. It provides valuable insights from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
One dataset versus multiple models refers to the case when a single dataset is used to train multiple models. Table 6 shows the top 10 datasets based on the number of models trained on them. In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks. It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets. We observed that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets. Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
Table 6. Top-10 datasets sorted by # of models trained.
Dataset Total Fine-tune Adapter Quantization Merges